OpenStudy (anonymous):

At a new job, you have earned $1,800, $2,100, $1,600, $1,900, and $2,300 the first 5 months. What do you need to earn in the sixth month to have an average of $2,000 for the first 6 months? A $2,000 B $2,100 C $2,200 D $2,300

OpenStudy (texaschic101):

(1800 + 2100 + 1600 + 1900 + 2300 + x)/6 = 2000 can you find it now ?

OpenStudy (texaschic101):

(1800 + 2100 + 1600 + 1900 + 2300 + x)/6 = 2000 (9700 + x)/6 = 2000 --- multiply both sides by 6 9700 + x = 12000 --- subtract 9700 from both sides x = 12000 - 9700 x = 2300 (D) any questions ?